OpenStudy (anonymous):

Solve this system for equations using substitution. Put your answer in ordered pair form(x,y) x=4y 3x-2y=20

OpenStudy (anas.p):

Place 4y in place of x in the second equation.... like \[3(4y)-2y=20\] Now you try.

OpenStudy (anonymous):

x=4y 3x-2y=20 3(4y)-2y=20 12y-2y=20 10y = 20 y= 2 put y = 2 in x= 4y x = 4*2 x = 8 (8,2)